### v1

@@ 3 Tereshchenkivs’ka Street, Kyiv-4, Ukraine, 01024
3d quantum mechanical systems with position dependent masses (PDM) admitting at least one second order integral of motion and symmetries with respect to dilatation or shift transformations are classified. Twenty seven such systems are specified and the completeness of the classification results is proved.
@@ 1 Introduction
One more interesting research field is formed by superintegrable systems with spin. The first example of such systems presented in paper [14], the systematic study of them started with papers [15, 16] where the systems with spin-orbit interaction were classified. Superintegrable systems with Pauli type interactions were studied in [17, 18] and [20], the examples of relativistic systems can be found in [20] and [21].
@@
It was shown in [24] the PDM Schr¨odinger equation can admit one, two, three, four, or six parametric Lie symmetry groups. In addition, there are also such equations which have no Lie symmetry. In other words, there are six well defined classes of such equations which admit n-parametric Lie groups with $n=1,2,3,4,6$ or do not have any Lie symmetry. And it is a natural idea to search for second order integrals of motion consequently for all these classes.
@@
The systems admitting two-parametric Lie groups and second order integrals of motion have been classified in [45]. We again find a number of new systems in addition to the known maximally superintegrable ones.
The natural next step is to classify the systems which admit one parametric Lie groups and second order integrals of motion. As it is shown in [24], up to equivalence these groups are reduced to dilatations, shifts along the fixed coordinate axis, rotations around this axis and some specific combinations of the mentioned transformations.
The integrable and superintegrable PDM admitting the symmetry w.r.t. the mentioned rotations, i.e., possessing the cylindric symmetry had been classified in paper [46]. The number of such systems appears to be rather extended. Namely, the 68 inequivalent systems admitting second order integrals of motion has been classified in [46]. The majority of these systems are neither maximally integrable nor degenerate and so are new and cannot be discovered using the approaches exploited in [28, 29, 30], [35, 36] and [37].
In the present paper we make the next step to the complete classification of the integrable and superintegrable PDM systems admitting one parametric Lie symmetry groups. Namely, we present the complete classification of two subclasses of such systems. The first class include the systems possessing the dilatation symmetry while the system belonging to the second class are invariant w.r.t. the shifts along the third coordinate axis.
@@
To classify the mentioned systems we use the optimised algorithm of solution of the related determining equations proposed in [46]. The number of the inequivalent systems with the dilatation symmetry appears to be equal to nine, including four integrable, three superintegrable and two maximally superintegrable ones. Notice that a special restricted class of such systems was studied earlier in papers [42] and [47].
@@
2 Main definitions
@@
$H=\frac{1}{4}(m^{\alpha}p_{a}m^{\beta}p_{a}m^{\gamma}+m^{\gamma}p_{a}m^{\beta}p_{a}m^{\alpha})+\hat{V}(x)$ (2)
where $p_{a}=-i\partial_{a}$, $m=m({\bf x})$ is a function of spatial variables ${\bf x}=(x_{1},x_{2},x_{3})$, associated with the position dependent mass, $\alpha,\beta$ and $\gamma$ are the so called ambiguity parameters satisfying the condition $\alpha+\beta+\gamma=-1$, and and summation from 1 to 3 is imposed over the repeating index $a$.
There are various physical speculations how to fix the ambiguity parameters in the particular models based on equation (2) . However, the systems with different values of these parameters are mathematically equivalent up to redefinition of potentials $V({\bf x})$. To simplify the following calculations we will fix them in the following manner: $\beta=0,\gamma=\alpha=-\frac{1}{2}$ and denote $f=f({\bf x})=\frac{1}{2m({\bf x})}$. As a result hamiltonian (2) is reduced to the following form
$H=f^{\frac{1}{2}}p_{a}p_{a}f^{\frac{1}{2}}+V({\bf x}).$ (3)
Alternatively, there are some historical reasons to consider the version $\alpha=\gamma=0,\beta=-1$ which corresponds to the following form of hamiltonian (2):
$H=p_{a}fp_{a}+\hat{V}.$ (4)
$V=\hat{V}+V^{k}$ (5)
where
and $\Delta$ is the Laplace operator.
Formula (6) represents an example of kinematical potentials which can be reduced to zero by the rearranging the ambiguity parameters.
In paper [24] all equations (1) with Hamiltonians (4) admitting at least one parametric Lie symmetry group were classified. It was shown that there are six inequivalent symmetries which can be accepted by the PDM Schrödinger equations. They include rotation around the third coordinate axis, shift along this axis and dilatation groups. In addition, we can fix three combined symmetries which are superpositions of the mentioned ones.
In paper [46] we find all inequivalent PDM systems invariant with respect to the mentioned rotations, which admit second order integrals of motion, i.e., differential operators of the following generic form
@@ 3 Determining equations
$\mu^{ab}V_{b}-f\eta_{a}+F^{a}=0$ (13)
where $\delta^{bc}$ is the Kronecker delta, $f_{n}=\frac{\partial f}{\partial x_{n}},\ \mu^{an}_{n}=\frac{\partial\mu^{an}}{\partial x_{n}}$, etc., and summation is imposed over the repeating indices $n$ over the values $n=1,2,3$. Moreover,
$F^{a}=(f\mu^{ak}_{km}-\mu^{mn}f_{na})_{m}$ (14)
for the Hamiltonian of form (4) and
$F^{a}=(f\mu^{ak}_{km}-\mu^{mn}f_{na})_{m}+\mu^{ab}V^{(k)}_{b}$ (15)
for the Hamiltonian of form (5).
The term (14) in general is not trivial. However, it generates a kinematical part of potential. On the other hand the term (14) is reduced to zero provided equations (11) and (12) are satisfied. And this is why it is reasonable to use just representation (5) for the Hamiltinian, since in this case equation (13) is reduced to the maximally simple form, i.e.,
$\mu^{ab}V_{b}-f\eta_{a}=0.$ (16)
@@
with arbitrary function $g({\bf x})$. In addition the conformal Killing tensors include fourth order polynomials in $\bf x$.
The determining equations (12) represent the coupled system of three nonlinear partial differential equation equations for two unknowns $g({\bf x})$ and $f({\bf x})$. However this system can be linearizing by introduction the new dependent variables
$M=\frac{1}{f},\ N=\frac{g}{f}$ (18)
which reduces (12) to the following form:
@@
Equation (16) in its turn can be effectively linearised by introducing the following new dependent variables $\tilde{M}$ and $R$:
$\tilde{M}=MV,\ R=\frac{N\tilde{M}}{M}-\eta$ (20)
which reduce (16) to the following equation:
$\left(\mu^{nn}_{a}+2\mu^{na}_{n}\right)\tilde{M}+5(\mu^{an}\tilde{M}_{n}+R_{a})=0$ (21)
@@
Notice that the related inverse mass potential can be expressed via $M$ and $\tilde{M}$ in the following form:
$f=\frac{1}{M},\ V=\frac{\tilde{M}}{M}$ (22)
while the corresponding functions $g$ and $\eta$ are expressed via $M$ and $\tilde{M}$ in the following manner:
$g=\frac{N}{M},\ \eta=-\frac{N\tilde{M}}{M}-R.$ (23)
Just the linearised determining equation (18) will be used in the following to solve our classification problem.
@@ 4.2 Decoupling of the determining equations
The autonomous subsystem (11) defines the conformal Killing tensor which is a linear combination of the term presented in (17) and the following terms (see, e.g., [48]) :
$\mu^{ab}_{1}=\lambda_{1}^{ab},$ (26)
$\mu^{ab}_{2}=\lambda_{2}^{a}x_{b}+\lambda_{2}^{b}x_{a}-2\delta^{ab}\lambda_{2}^{c}x_{c},$ (27)
$\mu^{ab}_{5}=\delta^{ab}r^{2}+k(x_{a}x_{b}-\delta^{ab}r^{2}),$ (30)
$\mu^{ab}_{6}=\lambda_{6}^{ab}r^{2}-(x_{a}\lambda_{6}^{bc}+x_{b}\lambda_{6}^{ac})x_{c}-\delta^{ab}\lambda_{6}^{cd}x_{c}x_{d},$ (31)
$\mu^{ab}_{7}=(x_{a}\lambda_{7}^{b}+x_{b}\lambda_{7}^{a})r^{2}-4x_{a}x_{b}\lambda_{7}^{c}x_{c}+\delta^{ab}\lambda_{7}^{c}x_{c}r^{2},$ (32)
$\mu^{ab}_{9}=\lambda_{9}^{ab}r^{4}-2(x_{a}\lambda_{9}^{bc}+x_{b}\lambda_{9}^{ac})x_{c}r^{2}+(4x_{a}x_{b}+\delta^{ab}r^{2})\lambda_{9}^{cd}x_{c}x_{d}+\delta^{ab}\lambda_{9}^{cd}x_{c}x_{d}r^{2}$ (34)
where $r=\sqrt{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}}$, $\lambda_{n}^{a}$ and $\lambda_{n}^{ab}$ are arbitrary parameters, satisfying the conditions $\lambda_{n}^{ab}=\lambda_{n}^{ba},\lambda_{n}^{bb}=0$. Thus we can represent the generic integral of motion (7) in the following form:
$Q=Q_{(0)}+\sum Q_{(n)}+\eta$ (35)
where
[…]
Formulae (17) and (26)-(34) include an arbitrary function and 35 arbitrary parameters, so the system of determining equations (19) is rather complicated. Happily the considered PDM systems by definition should be invariant w.r.t. the scaling transformations. It means that the related equations (19) cannot include linear combinations of all polynomials listed in (26)-(34) but only homogeneous ones. Indeed, integral of motion (7) including inhomogeneous terms by construction is not invariant w.r.t. the mentioned transformations and so can be transformed to linear combination of integrals whose coefficients are homogeneous polynomials with arbitrary multipliers. In other words, the determining equations (19) are reduced to the five decoupled subsystems corresponding to the Killing tensors which are $n$-order homogeneous polynomials with $n=0,1,2,3,4$. Moreover, since our Hamiltonians (3) with arbitrary elements (8) are invariant with respect to the inverse transformation (25) we can restrict ourselves to the polynomials of order $n<3$, since symmetries with $n$=3 and $n$=4 appears to be equivalent to ones with $n=1$ and $n=0$ correspondingly. The corresponding Killing tensors are $\mu_{0}^{ab}+\mu_{1}^{ab},\mu_{0}^{ab}+\mu_{2}^{ab}+\mu_{3}^{ab}$ and $\mu_{0}^{ab}+\mu_{5}^{ab}+\mu_{6}^{ab}$ which we represent in (17) and the following formulae:
@@
Notice that functions $g({\bf x})$ should be homogeneous functions of independent variables. Moreover their homogeneity degree should be zero in case (37), one in case (38) and two in case (39).
To classify the scale invariant PDM systems admitting second order integrals of motion it is sufficient to solve the determining equations (18) where $\mu^{ab}$ are special Killing tensors fixed in (38), (39). In other words, our classification problem is decoupled to three subproblems corresponding to symmetries whose differential terms are independent on $x_{a}$ or are linear and quadratic in these variables.

### v3

@@ 3 Tereshchenkivs’ka Street, Kyiv-4, Ukraine, 01024
3d quantum mechanical systems with position dependent masses (PDM) admitting at least one second order integral of motion and symmetries with respect to dilatation or shift transformations are classified. Twenty seven such systems are specified and the completeness of the classification results is proved. In this way the next step to the complete classification of integrable PDM system is realized.
@@ 1 Introduction
An interesting research field is formed by superintegrable systems with spin. The first example of such systems presented in paper [14], the systematic study of them started with papers [15, 16] where the systems with spin-orbit interaction were classified. Superintegrable systems with Pauli type interactions were studied in [17, 18] and [20], the examples of relativistic systems can be found in [20] and [21].
@@
It was shown in [24] the PDM Schrödinger equation can admit one, two, three, four, or six parametric Lie symmetry groups. In addition, there are also such equations which have no Lie symmetry. In other words, there are six well defined classes of such equations which admit n-parametric Lie groups with $n=1,2,3,4,6$ or do not have any Lie symmetry. And it is a natural idea to search for second order integrals of motion consequently for all these classes.
@@
The systems admitting two-parametric Lie groups and second order integrals of motion had been classified in [45]. We again find a number of new systems in addition to the known maximally superintegrable ones.
The natural next step is to classify the systems which admit one parametric Lie groups and second order integrals of motion. As it was shown in [24], up to equivalence there are six one parametric Lie groups which can be accepted by equation (1). These groups are:
  • Rotations around the fixed axis;
  • Dilatations;
  • Shifts along the fixed axis;
  • Superposition of the rotations and shifts;
  • superposition of the rotations and dilatation;
  • Supeposition of the rotations and the special conformal transformations.
The integrable and superintegrable PDM admitting the symmetry w.r.t. the mentioned rotations, i.e., possessing the cylindric symmetry had been classified in paper [46]. The number of such systems appears to be rather extended. Namely, the 68 inequivalent systems admitting second order integrals of motion have been discovered in [46]. The majority of these systems are neither maximally integrable nor degenerate and so are new and cannot be discovered using the approaches exploited in [28, 29, 30], [35, 36] and [37].
In the present paper we make the next step to the complete classification of the integrable and superintegrable 3d PDM systems admitting one parametric Lie symmetry groups. Namely, we present the complete classification of two subclasses of such systems. The first class includes the systems possessing the dilatation symmetry while the system belonging to the second class are invariant w.r.t. the shifts along the third coordinate axis. Just these symmetries are presented in Items 3 and 4. Formally speaking the results of paper [46] and of the present paper cover the half part of the possible one parametric symmetries enumerated in the above presented items. But in fact these results include the main part of the systems admitting these symmetries. The number of the systems presented in any of these papers is more extended that the total number of the systems admitting the symmetries enumerated in the three last items.
@@
To classify the mentioned systems we use the optimised algorithm of solution of the related determining equations proposed in [46], see Section 2 in the following text.
The number of the inequivalent systems with the dilatation symmetry appears to be equal to ten, including five integrable, three superintegrable and two maximally superintegrable ones. Notice that a special restricted class of such systems was studied earlier in papers [42] and [47].
@@
The paper is organized in the following manner. In addition to Introduction and Discussion sections it includes three large parts named as Sections 2, 3 and 4.
In Sections 2 we present the general discussion which includes the basic definitions (Section 2.1), the deduction of the determining equations (Section 2.2) and their optimization (Section 2.3). In Section 2.4 we formulate the algorithm for the construction of general solutions of the determining equations which is used in the present and our previous papers.
Rather extended Sections 3 and 4 include the solutions of the classification problems formulated in the above.
In Section 3 we classify the PDM systems which are invariant with respect to the dilatation transformations and admit second order integrals of motion. In subsection 3.1 we show that the related determining equations can be essentially simplified and decoupled to the subsystems which can be exactly solved. In subsections 3.2-3.4 we find and present the solutions of these subsystems together with the corresponding integrals of motion.
In Section 4 the shift invariant PDM systems are classified. In the beginning of this section we discuss the specific properties of the related integrals of motion and the ways to use these properties to optimise the solution procedure. The simplified determining equations are solved in subsection 4.1 which makes it possible to classify the systems admitting at least one second order integral of motion. The cases when the related systems are superintegrable and maximally superintegrable are specified in Subsection 4.2.
Both Sections 3 and 4 are ended by the classification tables when the obtained results are summarized. The last Section 5 includes the discussion of the obtained results.
2 Formulation of the problem and algorithm for its solution
2.1 Basic definitions
@@
$H=\frac{1}{2}(M^{\alpha}p_{a}M^{\beta}p_{a}M^{\gamma}+M^{\gamma}p_{a}M^{\beta}p_{a}M^{\alpha})+\hat{V}(x)$ (2)
where $p_{a}=-i\partial_{a}$, $M=M({\bf x})$ is a function of spatial variables ${\bf x}=(x_{1},x_{2},x_{3})$, associated with the position dependent mass, $\alpha,\beta$ and $\gamma$ are the so called ambiguity parameters satisfying the condition $\alpha+\beta+\gamma=-1$, and and summation from 1 to 3 is imposed over the repeating index $a$ (i.e., Einstein’s summation convention is used). We will use the summation convention for all repeating indices not necessary for one up and second down.
There are various physical speculations how to fix the ambiguity parameters in the particular models based on equation (2) . However, the systems with different values of these parameters are mathematically equivalent up to redefinition of potentials $V({\bf x})$. To obtain the most simple and compact form of the related potentials we will fix these parameters in the following manner: $\beta=0,\gamma=\alpha=-\frac{1}{2}$ and denote $f=f({\bf x})=\frac{1}{M({\bf x})}$. As a result hamiltonian (2) is reduced to the following form
$H=f^{\frac{1}{2}}p_{a}p_{a}f^{\frac{1}{2}}+V({\bf x})$ (3)
where $p_{a}p_{a}=p_{1}^{2}+p_{2}^{2}+p_{3}^{2}$. In paper [46] we find all inequivalent PDM systems invariant with respect to the mentioned rotations, which admit second order integrals of motion, i.e., differential operators of the following generic form
@@ 2.2 Determining equations
$\mu^{an}V_{n}-f\eta_{a}=0$ (10)
where $\delta^{bc}$ is the Kronecker delta, $f_{n}=\frac{\partial f}{\partial x_{n}},\ \mu^{an}_{n}=\frac{\partial\mu^{an}}{\partial x_{n}}$, etc., and summation is imposed over the repeating indices $n$ over the values $n=1,2,3$, i.e., $\mu^{an}f_{n}=\mu^{a1}f_{1}+\mu^{a2}f_{2}+\mu^{a3}f_{3}$ and $\mu^{an}V_{n}=\mu^{a1}V_{1}+\mu^{a2}V_{2}+\mu^{a3}V_{3}$.
2.3 Linearization of the determining equations
@@
with arbitrary function $g({\bf x})$. In addition the conformal Killing tensors include fourth order polynomials in $\bf x$ which are given by the following formulae: of the following form (see, e.g., [48]) :
$\mu^{ab}_{1}=\lambda_{1}^{ab},$ (12)
$\mu^{ab}_{2}=\lambda_{2}^{a}x_{b}+\lambda_{2}^{b}x_{a},$ (13)
$\mu^{ab}_{5}=kx_{a}x_{b},$ (16)
$\mu^{ab}_{6}=\lambda_{6}^{ab}r^{2}-(x_{a}\lambda_{6}^{bc}+x_{b}\lambda_{6}^{ac})x_{c},$ (17)
$\mu^{ab}_{7}=(x_{a}\lambda_{7}^{b}+x_{b}\lambda_{7}^{a})r^{2}-4x_{a}x_{b}\lambda_{7}^{c}x_{c},$ (18)
$\mu^{ab}_{9}=\lambda_{9}^{ab}r^{4}-2(x_{a}\lambda_{9}^{bc}+x_{b}\lambda_{9}^{ac})x_{c}r^{2}+(4x_{a}x_{b}+\delta^{ab}r^{2})\lambda_{9}^{cd}x_{c}x_{d}$ (20)
where $r=\sqrt{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}}$, $\lambda_{n}^{a}$ and $\lambda_{n}^{ab}$ are arbitrary parameters, satisfying the conditions $\lambda_{n}^{ab}=\lambda_{n}^{ba},\lambda_{n}^{bb}=0$.
Thus we can represent the generic integral of motion (4) in the following form:
$Q=Q_{(0)}+\sum Q_{(n)}+\eta$ (21)
where
[…]
We have in hands the generic solution of the first subsystem of the determining equations presented in (8). The remaining determining equations, i.e., (9) and (10) represent the coupled system of three nonlinear partial differential equation equations for two unknowns $g({\bf x})$ and $f({\bf x})$. However this systems can be linearized by introducing the new dependent variables $M$ and $N$ connected with $f$ and $g$ in the following manner:
$f=\frac{1}{M},\ g=\frac{V}{M}.$ (23)
As a result we transform subsystem (9) to the following form:
@@
Equation (10) in its turn can be effectively simplified by introducing the new dependent variables $\tilde{M}$ and $\tilde{N}$ in the following way:
$V=\frac{\tilde{M}}{M},\ \eta=\frac{N\tilde{M}}{M}-\tilde{N}$ (25)
which reduce (10) to the following form:
$\left(\mu^{nn}_{a}+2\mu^{na}_{n}\right)\tilde{M}+5(\mu^{an}\tilde{M}_{n}+\tilde{N}_{a})=0$ (26)
@@
Just the linearised determining equation (24) will be used in the following to solve our classification problem. Solving this equation we will find the mass functions of the PDM systems admitting second order integrals of motion of generic form (21). In fact in this way we will immediately find the corresponding potential also since equation (26) coincides with (24).
2.4 The algorithm for searching of general solution of the determining equations
Equation (24) looks rather gentle but in fact we are supposed to deal with a very complicated system of the coupled partial differential equations, including three dependent and three independent variables. This complexity is caused by the fact that the conformal Killing tensor $\mu^{ab}$ present in (24) includes as many as 35 arbitrary parameters, refer to (12)-(20). However, for the case of the PDM systems admitting at least one parametric Lie symmetry groups this system can be effectively simplified and decoupled to the subsystems with the essentially reduced number of these parameters.
The algorithm we use to classify the PDM systems admitting second order integrals of motion and one parametric Lie groups includes the following steps.
1. To deduce the determining equations for the variable coefficients of the integrals of motion. This step was represented in the beginning of Section 2.
2. To linearize the determining equations. It has been done in the previous section where the linear versions of the determining equations were obtained, refer to equations (24) and (26).
3. To solve step by step equations (24) for all the cases when the PDM system admits one of the symmetry groups enumerated in the items present in the Introduction. Notice that in any of the mentioned cases the determining equations are much more simple than in the generic case since the arbitrary elements $f$ and $V$ are essentially restricted by the symmetry condition. For example, in the case of the complete rotation invariance of the considered PDM system these arbitrary elements should be functions of the only (radial) variable, and the determining equations are reduced to the ordinary differential ones [42].
4. In addition to the previous note, for the cases when a Lie symmetry is present the systems of the determining equations can be effectively decoupled to the relatively simple subsystems which can be solved exactly. Just such decoupling is the next step of our algorithm.
The ways to the mentioned decoupling are strongly dependent on the postulated Lie symmetries. In the case of symmetry with respect to rotations the decoupled subsystems describe separately the scalar, vector and tensor integrals of motion with different fixed parities [46]. There are also specific ways to the decoupling for the cases of dilatation and shifts invariance which are presented in the following text, see Section 3.2 and the beginning of Section 4
5. The effective tools for the additional simplification of the determining equations are presented by the continuous and discrete equivalence transformation. For the systems analysed in the present paper such transformations are discussed in Section 2.3.
6. The last step of the algorithm is to find the generic solutions of the decoupled and simplified subsystems taking the care on the possible presence of special solutions. It is important to stress that this step is by no means trivial or easy since the simplified subsystems are still the systems of coupled partial differential equations with variable coefficients. However, using special and some times rather sophisticated approaches we were able to find the related exact general solutions in explicit form.
In the following sections we apply steps 4-6 of the presented algorithm for the classification of PDM systems admitting second order integrals of motion and one parametric Lie groups specified in Items 2 and 3 presented in Introduction, i.e, groups of dilatation and shift transformations. The cases of the remaining groups specified in Items 4-6 will be considered in the following (and final) paper of this series.
@@ 3.2 Decoupling of the determining equations
Formulae (11) and (12)-(20) include an arbitrary function and 35 arbitrary parameters, so the system of determining equations (24) is rather complicated. Happily the considered PDM systems by definition should be invariant w.r.t. the scaling transformations. It means that the related equations (24) cannot include linear combinations of all polynomials listed in (12)-(20) but only homogeneous ones. Indeed, integral of motion (4) including inhomogeneous terms by construction is not invariant w.r.t. the mentioned transformations and so can be transformed to linear combination of integrals whose coefficients are homogeneous polynomials with arbitrary multipliers. In other words, the determining equations (24) are reduced to the five decoupled subsystems corresponding to the Killing tensors which are $n$-order homogeneous polynomials with $n=0,1,2,3,4$. Moreover, since our Hamiltonians (3) with arbitrary elements (5) are invariant with respect to the inversion transformation (28) we can restrict ourselves to the polynomials of order $n<3$, since symmetries with $n$=3 and $n$=4 appears to be equivalent to ones with $n=1$ and $n=0$ correspondingly. The corresponding Killing tensors are $\mu_{0}^{ab}+\mu_{1}^{ab},\mu_{0}^{ab}+\mu_{2}^{ab}+\mu_{3}^{ab}$ and $\mu_{0}^{ab}+\mu_{5}^{ab}+\mu_{6}^{ab}$ which we represent in (11) and the following formulae:
@@
Notice that $g({\bf r})$ should be homogeneous functions of independent variables. Moreover their homogeneity degree should be zero in case (29), one in case (30) and two in case (31).
To classify the scale invariant PDM systems admitting second order integrals of motion it is sufficient to solve the determining equations (23) where $\mu^{ab}$ are special Killing tensors fixed in (30), (31). In other words, our classification problem is decoupled to three subproblems corresponding to symmetries whose differential terms are independent on $x_{a}$, are linear in $x_{a}$ or are quadratic in these variables.